OpenStudy (anonymous):

HELP 2x^2+8x-64 / x-8 * 6x/3x^2-12x

OpenStudy (dan815):

hi

OpenStudy (dan815):

can u factor some stuff

OpenStudy (anonymous):

idk how to . ive been havin help.

OpenStudy (dan815):

ah u want to see video on it this whole problem is just factoring and pulling common variables out and simplifying

OpenStudy (dan815):

lets do the easy one for now 6x/3X^2-12x =6x/3x(x-4) =2/(x-4) thats the simplified form of that 2nd fraction now learn factoring and do the first part of the fraction :)

OpenStudy (anonymous):

i guess so .

OpenStudy (anonymous):

wtf ? you just confused me even more. lmao

OpenStudy (dan815):

http://www.youtube.com/watch?v=qy0uSglDxww

OpenStudy (dan815):

wut i just pulled 3 and x was common -_-

OpenStudy (anonymous):

i get confused really easy, as you can tell .

OpenStudy (dan815):



OpenStudy (dan815):

this is the 2nd fraction of ur question, for the first part u need to learn factoring

OpenStudy (anonymous):

okay, -_- mean

OpenStudy (dan815):

u know how to do that tho right?? im just pulling out the common factors

OpenStudy (dan815):

and simplifying